OpenStudy (anonymous):

Find the derivative: Y=x^3(5x^2+1)^-2/3

OpenStudy (aravindg):

help me plzzz

OpenStudy (anonymous):

the answer is 1/3x^2(5x^2+1)^-5/3(25x^2+9) BUT HOW?

OpenStudy (anonymous):

3x^2(5x^2+1)^-2/3 + -2/3(5x^2+1)^-5/3(10x)x^3

OpenStudy (anonymous):

not sure what you did there. did you use chain rule first somehow? my first step was using the product rule.

OpenStudy (mimi_x3):

  Use the product rule u = x^3 u' = 3x^2 v = (5x^2+1)^-2/3 v' = -20x / 3(5x^2+1)^3 = 1/3(5x^2+1)^2